# Quasi-Clique Discovery via Energy Diffusion

### v2

@@
Abstract
Discovering quasi-cliques—subgraphs whose edge density exceeds a given threshold—is a fundamental task in graph mining with applications to web spam detection, fraud screening, and e-commerce recommendation. However, existing methods for quasi-clique discovery on large-scale web graphs are often sensitive to random seeds or lack of explicit edge-density guarantees, making the task challenging in practice. This paper presents EDQC, an energy diffusion–based method for quasi-clique discovery. EDQC first employs an adaptive energy diffusion process to generate an energy ranking that highlights structurally cohesive regions. Guided by this energy ranking, the algorithm identifies a high-quality subgraph by minimizing conductance, a standard measure from community detection. This subgraph is then refined to meet the specified density threshold. Extensive experiments on 75 real-world graphs show that EDQC finds larger quasi-cliques on most datasets, with consistently lower variance across runs and competitive runtime. To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into quasi-clique discovery.
1 Introduction
Dense subgraph models are fundamental to graph mining on large-scale web graphs, enabling web spam detection gibson2005discovering ; lin2021mining , fraud screening zheng2019author ; fang2022densest , and e-commerce recommendation schafer2001commerce ; song2021subgraph . Among these models, the clique represents the most stringent form, requiring that every pair of vertices be directly connected to each other saito2007large . However, due to noise and missing links in real-world graphs, exact cliques are rarely observed. To relax this strict requirement, several approximate dense subgraph models have been proposed, including quasi-cliques abello2002massive ; pei2005mining , $k$-clubs mokken1979cliques , and $k$-plexes seidman1978graph . In particular, the $\gamma$-quasi-clique, defined by a minimum edge-density threshold, is one of the most widely adopted formulations chen2021nuqclq ; pang2024similarity ; liu2024optimization .
Given an undirected graph $G$ and an edge-density threshold $\gamma\in(0,1]$, the Maximum Quasi-Clique Problem (MQCP) seeks the largest vertex set $S$ such that the induced subgraph $G[S]$ exhibits an edge density of at least $\gamma$. This edge-density criterion provides a more direct measure of structural cohesion compared to the degree-based definitions utilized by methods such as FastQC konar2024optimal and IterQC liu2008effective . However, the MQCP is NP-hard even for fixed values of $\gamma$ pattillo2013maximum and does not possess the hereditary property that allows for efficient pruning in problems related to cliques xia2025maximum . Nonetheless, quasi-clique discovery remains highly valuable, as real-world graphs typically exhibit dense yet incomplete connectivity rather than perfect cliques, making quasi-cliques crucial for identifying significant or anomalous substructures.
Figure 1: A clique and a 0.8-quasi-clique.
One practical application of quasi-clique discovery is identifying link farm spam pages in web graphs pang2024similarity , where densely connected spam pages form $\gamma$-quasi-cliques, enabling the identification of suspicious link farms even under noise and missing links. As shown in Figure 1, the dashed edges in the left subgraph form a size-4 clique, representing a fully connected link farm. In the right subgraph, the dashed edges form a 0.8-quasi-clique, still suspicious despite one missing link, capturing a potential link farm under partial observation or evasion.
Despite recent progress, discovering quasi-cliques in real large-scale web graphs remains challenging due to both scale and combinatorial complexity. While exact algorithms for MQCP, typically based on integer programming or techniques such as branch-and-bound uno2010efficient ; pastukhov2018maximum ; ribeiro2019exact ; rahman2024fast , guarantee optimal solutions, their exponential time complexity restricts their applicability to large graphs. As a practical alternative, numerous heuristic methods have been proposed to enable scalable quasi-clique discovery, which can be roughly grouped into four categories. First, greedy expansion techniques, such as the greedy strategy with local refinement introduced in tsourakakis2013denser and the restart-enabled search in RIG oliveira2013construction , prioritize speed but are vulnerable to becoming trapped in local optima. Second, metaheuristic frameworks like TSQC djeddi2019extension and OBMA zhou2020opposition employ tabu search and opposition-based learning, offering enhanced robustness at a higher computational cost. Third, local search algorithms, including NuQClq chen2021nuqclq and its variants NuQClqF1 and NuQClqR1 liu2024optimization , utilize bounded configuration checking and adaptive rules for candidate refinement. However, despite their sophisticated local optimizations, these methods often exhibit instability due to their dependence on random seeds. Lastly, similarity-based heuristics such as NBSim and its accelerated variant, FastNBSim, utilize vertex neighborhood overlap to identify and expand quasi-cliques pang2024similarity . While both methods are highly efficient, FastNBSim introduces MinHash-based approximations, making it sensitive to random seeds and potentially less stable. Moreover, neither method supports a user-specified edge-density threshold. Instead, the edge-density of the returned subgraph emerges implicitly from similarity parameters. Consequently, there is no guarantee that the output meets a specified minimum edge-density requirement. This limits their suitability for applications that require strict control over quasi-clique edge-density thresholds.
Among these four categories, local search and similarity-based methods have been the most competitive for discovering large quasi-cliques. The state-of-the-art algorithms for MQCP include NBSim and its accelerated variant, FastNBSim pang2024similarity , as well as the local search algorithms NuQClqF1 and NuQClqR1 liu2024optimization . However, these methods remain sensitive to random seeds or lack of explicit edge-density guarantees.
Energy diffusion is a natural candidate for addressing the challenges above. Diffusion methods such as PageRank page1999pagerank and the heat kernel chung2007heat have shown strong performance on structure-aware tasks, e.g., vertex ranking and community detection kloster2014heat ; gleich2015multilinear , where diffused mass tends to localize within cohesive regions. These tasks overlap with quasi-clique discovery in highlighting tightly interlinked vertices, yet they are not equivalent because quasi-cliques must satisfy a specified minimum edge-density on induced subgraphs. This connection motivates exploring energy diffusion for quasi-clique discovery to reduce sensitivity to random seeds in practice and to furnish a stable candidate region for density-constrained extraction.
This paper introduces EDQC, an energy diffusion-based method for quasi-clique discovery. EDQC begins with an adaptive energy diffusion phase, where the energy transfer rate is dynamically adjusted based on the local energy distribution. This phase runs for a fixed number of iterations from each source vertex, generating an energy ranking that highlights structurally cohesive regions. Guided by this energy ranking, EDQC then employs a principled extraction mechanism from community detection theory. It applies a sweep cut to the energy-ranked vertices to identify a high-quality subgraph that minimizes conductance. This subgraph is subsequently refined through a structure-driven procedure to guarantee it meets the specified density threshold. Extensive experiments on 75 real-world graphs, including classic web graphs, demonstrate that EDQC discovers larger quasi-cliques on most datasets, with consistently lower variance and competitive runtime. To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into quasi-clique discovery.
2 Preliminaries
@@
All graphs considered in this paper are undirected and simple (no loops or parallel edges). Let $G=(V,E)$ represent a graph with a vertex set $V$ and an edge set $E$. Two vertices $u$ and $v$ within $V$ are termed adjacent if $uv\in E$, and one vertex is referred to as a neighbor of the other. The set of all neighbors for a vertex $u$ is known as its neighborhood, denoted $N(u)$, while the closed neighborhood of $u$ is denoted $N[u]=N(u)\cup\{u\}$. The degree of a vertex $u$, denoted as $d(u)$, indicates the number of its neighbors, i.e., $d(u)=|N(u)|$.
@@
Let $n=|V|$ and $m=|E|$. For a vertex set $S\subseteq V$, the induced subgraph is $G[S]$ with vertex set $S$ and edge set $E(S)=\{\{u,v\}\in E\mid u,v\in S\}$. Let $m(S)=|E(S)|$ denote the number of internal edges in $G[S]$, and define the internal degree of $u\in S$ by
$d_{S}(u)=|N(u)\cap S|.$
The edge density of $S$ (and of $G[S]$) is
$\eta(S)=\begin{cases}\dfrac{2\,m(S)}{|S|(|S|-1)},&|S|\geq 2,\\[6.0pt] 0,&otherwise,\end{cases}$
and $\eta(G)$ is shorthand for $\eta(V)$. Throughout this paper, unless otherwise stated, density means edge density.
@@
Given $G=(V,E)$ and a density threshold $\gamma\in(0,1]$. A vertex set $S\subseteq V$ induces a $\gamma$-quasi-clique if the induced subgraph $G[S]$ has edge density at least $\gamma$, i.e.,
$\eta(G[S])\;\geq\;\gamma.$
For brevity, we also refer to $G[S]$ itself as a $\gamma$-quasi-clique. A $\gamma$-quasi-clique $G[S]$ is maximal if there is no $v\in V\setminus S$ such that $\eta\!\left(G[S\cup\{v\}]\right)\geq\gamma$.
@@
The induced subgraph $G[S^{*}]$ is a maximum $\gamma$-quasi-clique. When $\gamma=1$, MQCP reduces to the classical Maximum Clique Problem. For any fixed $\gamma$, MQCP is NP-hard and may admit multiple optimal solutions pattillo2013maximum .
@@
Conductance is a standard measure used to evaluate how well a vertex set forms a community andersen2006local . For a non-empty, proper subset of vertices $S\subset V$, its volume is defined as $\mathrm{vol}(S)=\sum_{u\in S}d(u)$, and the cut between $S$ and its complement is $\mathrm{cut}(S,V\setminus S)=|\{\{u,v\}\in E\mid u\in S,v\in V\setminus S\}|$. The conductance of $S$ is then:
@@
A lower conductance value indicates a more clearly defined community, signifying few edges connecting the set to the rest of the graph relative to its internal connectivity.
3 The EDQC Algorithm
This section details the proposed Energy Diffusion-based Quasi-Clique discovery (EDQC) algorithm. Given a graph $G=(V,E)$ and a density threshold $\gamma$, EDQC identifies a large vertex set $S$ that induces a $\gamma$-quasi-clique. The overall pipeline is outlined in Algorithm 1.
The algorithm iterates through vertices, treating each as a source to initiate a two-stage process. First, an adaptive energy diffusion process generates an energy distribution over the graph. Second, guided by this energy distribution, a principled extraction mechanism identifies a high-quality quasi-clique. EDQC maintains the largest quasi-clique found across all source vertices. Processing vertices in non-increasing degree order is an efficiency heuristic rossi2014fast , as it ensures high-degree vertices are processed as energy sources first, thereby increasing the likelihood of finding a large quasi-clique early in the search.
@@ 6    $S^{*}\leftarrow S$;
3.1 Adaptive Energy Diffusion
The first stage, detailed in Algorithm 2, simulates the diffusion of energy from a source vertex $v$ over $T$ synchronous rounds. The goal is to produce an energy function $f:V\to\mathbb{R}_{\geq 0}$ where energy is concentrated in structurally cohesive regions surrounding the source.
The diffusion process is designed to be adaptive. The fraction of energy a vertex releases depends on its current energy relative to its neighborhood, a dynamic governed by an adaptive fraction $\alpha(u)$. This design is deliberate and distinct from standard diffusion mechanisms such as the heat kernel chung2007heat or PageRank page1999pagerank , which often treat connections uniformly. Our adaptive fraction makes the diffusion process sensitive to the evolving energy landscape, enabling EDQC to be more conservative in sparse regions and more aggressive within dense ones.
The diffusion starts by initializing a unit energy at the source vertex $v$ (line 1). In each of the $T$ synchronous rounds (lines 2–14), the algorithm first identifies the set of active vertices, $A$, based on the activation threshold $\theta$ (line 3). To ensure updates are synchronous, a temporary copy of the energy function, $f^{\prime}$, is created (line 4). The diffusion rule is then applied to each active vertex $u\in A$ with a non-zero degree ($d(u)>0$). Isolated vertices are skipped as they have no neighbors to which energy can be transferred, a condition necessary for the conservation of total energy. For each such vertex $u\in A$ (line 5), the algorithm computes its adaptive fraction $\alpha(u)$ and the total energy to transfer, $\delta(u)$ (lines 7–8). This energy is then stochastically allocated among its neighbors using weights $\phi_{uw}$ derived from an exponential distribution (lines 9–10). Subsequently, the energies of the active vertex and its neighbors are updated in the temporary function $f^{\prime}$ (lines 11–13). After all active vertices have been processed, the energy function $f$ is updated with the temporary copy, completing the round (line 14).
@@ 17 $f\leftarrow f^{\prime}$;
This diffusion process has several key properties.
@@
Consider a fixed round and an arbitrary active vertex $u$ where $d(u)>0$. The synchronous updates associated with $u$ results in a change of $\delta(u)$ for $u$ and an addition of $\delta(u)\phi_{uw}$ for each neighbor $w\in N(u)$, satisfying $\sum_{w\in N(u)}\phi_{uw}=1$. Therefore, $\delta(u)=\sum_{w\in N(u)}\delta(u)\phi_{uw}$. Summing over all active vertices leads to the conclusion that $\sum_{x\in V}f^{\prime}(x)=\sum_{x\in V}f(x)$. Here, we should note that the energy of vertices for which $d(u)=0$ or $f(u)=0$ remains unchanged in the energy diffusion process. ∎
The conservation of energy is a crucial property for the algorithm’s stability. It guarantees that the process only redistributes the initial energy, rather than arbitrarily creating or losing it, thus ensuring a fair basis for comparing the energy distributions originating from different source vertices.
@@
The adaptive nature of our diffusion is formally established by this proposition. The bounds on $\alpha(u)$ guarantee a process that is both consistently active and stable, preventing the diffusion from stalling. Concurrently, the monotonicity is the key mechanism driving the algorithm’s core premise: it mathematically ensures that energy is exchanged more readily within already high-energy (and thus likely dense) regions, accelerating the concentration of energy.
@@
The expected fraction of energy that an active vertex $u$ transfers to a specific neighbor $w\in N(u)$ is uniform across all neighbors: $\mathbb{E}[\phi_{uw}]=1/d(u)$.
@@
While the exponential sampling introduces stochasticity to explore different structural paths, this proposition guarantees that the process remains unbiased in expectation. On average, all neighbors are treated fairly, ensuring that the algorithm itself introduces no arbitrary structural bias during the diffusion.
@@
The EnergyDiffusion algorithm runs in $O(T(n+m))$ time.
@@
The algorithm performs $T$ rounds. We analyze the computational cost of a single round. First, identifying the active set $A$ requires a full scan of all vertices, which takes $O(n)$ time. Next, the algorithm iterates through each active vertex $u\in A$. For each such vertex, all calculations—including the summations for $\alpha(u)$ and $\phi_{uw}$, and the subsequent updates—are proportional to its degree, $d(u)$. Therefore, the total work for processing all active vertices in a round is bounded by the sum of degrees of all vertices in the graph, which is $\sum_{u\in V}O(d(u))=O(m)$. Summing the costs within a single round gives a complexity of $O(n+m)$. As the process repeats for $T$ rounds, the total complexity of the algorithm is $O(T(n+m))$. ∎
3.2 Quasi-Clique Extraction
The second stage of EDQC takes the final energy function $f$ returned by Algorithm 2 and produces a vertex set $S$ that induces a $\gamma$-quasi-clique. This extraction process (Algorithm 3) has three main operations: an initial ranking of vertices, the identification of a high-quality initial subgraph, and a refinement procedure.
First, the algorithm identifies the set of active vertices $A$ and sorts them in descending order of their energy scores to create an ordered sequence $P$ (lines 1–2). This ranking is crucial, as it places vertices most likely to belong to a cohesive community at the forefront.
Next, to identify a high-quality initial candidate set, the algorithm performs a sweep cut andersen2006local on the energy-ranked list $P$ (line 3). This procedure examines the nested prefixes of the list and selects the one that minimizes conductance, as it represents the most well-defined community structure. This strategy is highly effective because the initial energy ranking provides a high-quality vertex permutation. It is a known principle in local graph clustering that sweep cuts perform best when applied to an ordering that places vertices from the same community contiguously andersen2006local . Our energy diffusion process is designed for precisely this purpose: it creates an ordering where structurally cohesive vertices are naturally placed at the top. This, in turn, enables the sweep cut to accurately identify the boundary of the best candidate subgraph.
Finally, this initial set $S$ undergoes a refinement procedure (lines 4–15). The refinement begins with a pruning loop (lines 4–6). If the initial set $S$ does not meet the density threshold, this loop iteratively removes the vertex with the minimum internal degree $d_{S}(u)$ until the set $S$ induces a valid $\gamma$-quasi-clique. Following this, an expansion phase seeks to enlarge the set (lines 8–15). In each pass, the remaining candidates $R$ are re-sorted based on their connectivity to the current set $S$ (line 10). The algorithm then iterates through this sorted list $R$, attempting to greedily add the first vertex $v$ that maintains the density constraint (lines 12–13). Upon a successful insertion, the current iteration over the candidate list is immediately terminated (line 15), allowing the algorithm to restart the process with an up-to-date candidate ranking in the next iteration.

### v3

# Cohesive Group Discovery in Interaction Graphs under Explicit Density Constraints

@@
Abstract.
Discovering cohesive groups is a fundamental primitive in graph-based recommender systems, underpinning tasks such as social recommendation, bundle discovery, and community-aware modeling. In interaction graphs, cohesion is often modeled as the $\gamma$-quasi-clique, an induced subgraph whose internal edge density meets a user-defined threshold $\gamma$. This formulation provides explicit control over within-group connectivity while accommodating the sparsity inherent in real-world data. This paper presents EDQC, an effective framework for cohesive group discovery under explicit density constraints. EDQC leverages a lightweight energy diffusion process to rank vertices for localizing promising candidate regions. Guided by this ranking, the framework extracts and refines a candidate subgraph to ensure the output strictly satisfies the target density requirement. Extensive experiments on 75 real-world graphs across varying density thresholds demonstrate that EDQC identifies the largest mean $\gamma$-quasi-cliques in the vast majority of cases, achieving lower variance than the state-of-the-art methods while maintaining competitive runtime. Furthermore, statistical analysis confirms that EDQC significantly outperforms the baselines, underscoring its robustness and practical utility for cohesive group discovery in graph-based recommender systems.
Cohesive group discovery, quasi-clique, graph-based recommender systems
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM SIGIR Conference 2026; July 20–24, 2026; Melbourne — Naarm, Australiaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
1. Introduction
Cohesive group discovery plays a fundamental role in graph-based recommender systems, enabling tasks such as social recommendation (Alina Christensen and Schiaffino, 2014), bundle discovery (Nguyen et al., 2024), and community-aware modeling (Guan et al., 2021). In many recommender settings, user–item interactions or user–user relations are naturally modeled as interaction graphs, where vertices denote users or items and edges capture affinity signals such as co-consumption or behavioral similarity (Schafer et al., 2001; Zhou, 2025). This representation encompasses various network types, including social, communication, citation, collaboration, and web graphs. Within such graphs, cohesive groups are often instantiated as dense induced subgraphs, supporting candidate expansion, group-aware ranking, and coordinated abnormal behavior detection.
Among dense subgraph models, the clique is the most stringent, requiring every pair of vertices to be connected (Saito et al., 2007). However, in real-world networks, noise, missing edges, and incomplete connectivity make exact cliques rare. To address this, several relaxed models have been proposed, including quasi-cliques (Abello et al., 2002; Pei et al., 2005), $k$-clubs (Mokken et al., 1979), and $k$-plexes (Seidman and Foster, 1978). In particular, the edge-density-based formulation, often referred to as the $\gamma$-quasi-clique, is widely adopted for its intuitive and explicit notion of structural cohesion (Chen et al., 2021; Pang et al., 2024; Liu et al., 2024).
Given an undirected graph $G$ and an edge-density threshold $\gamma\in(0,1]$, the Maximum Quasi-Clique Problem (MQCP) asks for a largest vertex set $S$ such that the induced subgraph $G[S]$ has edge density at least $\gamma$. Throughout the paper, density refers to edge density, and quasi-clique refers to the $\gamma$-quasi-clique unless otherwise stated. Unlike degree-based criteria used in FastQC (Konar and Sidiropoulos, 2024) and IterQC (Liu and Wong, 2008), MQCP enforces an explicit, user-specified density threshold. It is also distinct from community detection, which typically optimizes a global objective (e.g., modularity or conductance) to partition or cover the entire graph. MQCP instead extracts a single dense local vertex set, aligning with settings where one seeks a single cohesive set of users or items rather than a full partition. In recommender applications, $\gamma$ serves as a minimum cohesion requirement that explicitly controls within-group connectivity. Hence, methods that do not enforce a user-specified $\gamma$ may produce subgraphs that fail to satisfy the target density, limiting their utility as reliable constraint-aware components for recommendation tasks.
MQCP is NP-hard even for a fixed $\gamma$ (Pattillo et al., 2013) and lacks the hereditary property used for pruning in clique-finding algorithms (Xia et al., 2025). In particular, a subset of a $\gamma$-quasi-clique is not necessarily a $\gamma$-quasi-clique. Despite these challenges, MQCP remains practically relevant, since $\gamma$-quasi-cliques capture compact interest groups and suspicious behavioral clusters in real-world graphs with dense but imperfect connectivity.
Figure 1 illustrates a typical situation in interaction graphs. In the left subgraph, the dashed edges form a size-4 clique, whereas in the right subgraph one edge is missing but the dashed edges still constitute a $0.8$-quasi-clique. In practice, interactions are often noisy or partially observed, so nearly-complete groups like this should still be treated as coherent interest groups. This motivates studying $\gamma$-quasi-cliques as an explicit density-constrained model of structural cohesion.
Figure 1. A size-4 clique and a 0.8-quasi-clique. Dashed edges indicate observed interaction affinities.
Despite recent progress, discovering quasi-cliques in large graphs remains challenging due to both scale and combinatorial complexity. Exact MQCP solvers based on integer programming and branch-and-bound (Uno, 2010; Pastukhov et al., 2018; Ribeiro and Riveaux, 2019; Rahman et al., 2024) provide optimality guarantees but are restricted to small graphs in practice. Consequently, heuristics are often preferred for real-world graphs, and existing methods can be roughly grouped into four categories.
First, greedy expansion methods grow candidates with lightweight refinement (Tsourakakis et al., 2013; Oliveira et al., 2013). While fast, they are prone to getting trapped in poor local optima. Second, metaheuristic frameworks such as TSQC and OBMA (Djeddi et al., 2019; Zhou et al., 2020) enhance exploration using strategies like tabu search or opposition-based learning, but typically incur higher computational overhead, which limits scalability on large graphs. Third, local search algorithms, notably NuQClq and its variants (Chen et al., 2021; Liu et al., 2024), iteratively refine candidates with adaptive move rules (e.g., bounded configuration checking); however, their performance can vary substantially across random seeds on large-scale instances. This seed sensitivity hinders reproducibility and yields unstable outputs across runs, which is undesirable in recommender systems. Fourth, similarity-based heuristics such as NBSim and FastNBSim (Pang et al., 2024) exploit neighborhood overlap for efficiency, but they do not take a user-specified edge-density threshold $\gamma$ as input. Instead, density is controlled indirectly via similarity parameters, so the output is not guaranteed to satisfy a target $\gamma$. As a result, their outputs may be dense under an internal similarity criterion yet still violate a required edge-density threshold, making them unsuitable when $\gamma$ is a hard constraint. Among these categories, local search and similarity-based heuristics are often among the strongest baselines for mining large quasi-cliques. In particular, representative methods include NBSim/FastNBSim (Pang et al., 2024) and the local search algorithms NuQClqF1/NuQClqR1 (Liu et al., 2024). However, these methods either exhibit sensitivity to random seeds, or do not take a user-specified density threshold $\gamma$ as an explicit constraint and thus cannot guarantee meeting a target $\gamma$.
Energy diffusion provides a natural mechanism to improve the stability of quasi-clique discovery. Diffusion-based techniques such as PageRank (Page et al., 1999) and the heat kernel (Chung, 2007) have been effective for structure-aware tasks, including vertex ranking and local community detection (Kloster and Gleich, 2014; Gleich et al., 2015). A simple intuition is that short-range propagation tends to retain more mass within well-connected regions due to repeated exchanges along internal edges, which can be used to produce a locality-aware ranking from a source vertex. However, classical diffusion models are designed to approximate a probability distribution and are not tailored to extracting a vertex set that explicitly satisfies a density threshold $\gamma$.
To bridge this gap, we present EDQC, an effective framework that leverages energy diffusion for quasi-clique discovery. By utilizing lightweight primitives and sidestepping the iterative overhead often associated with metaheuristic search, EDQC is designed to be practical across different graph sizes. EDQC first runs an adaptive energy diffusion process to obtain an energy ranking, where the release rate of each vertex is adjusted by the energy in its neighborhood to concentrate scores within locally cohesive regions. While the ranking highlights promising vertices, it does not define a subgraph boundary. We therefore apply a conductance-based sweep cut to extract a well-separated candidate region, followed by a refinement step that prunes low-degree vertices and greedily adds qualifying vertices. The refinement enforces feasibility, ensuring that the output meets the user-specified density threshold $\gamma$.
Our main contributions are summarized as follows:
  • To our knowledge, we are the first to integrate energy diffusion into a heuristic framework for MQCP, using adaptive diffusion-based scoring to localize cohesive regions in interaction graphs at scale.
  • We design a diffusion-guided framework that converts local rankings into feasible $\gamma$-quasi-cliques through conductance-based extraction and density-constrained refinement, ensuring that the output satisfies the user-specified density threshold $\gamma$.
  • We evaluate EDQC on 75 real-world graphs with ablation studies and statistical significance tests, showing that it finds larger quasi-cliques than strong baselines on most datasets, with lower variance across runs and competitive runtime.
2. Preliminaries
@@
All graphs are undirected and simple (no loops or parallel edges). Let $G=(V,E)$ denote a graph with vertex set $V$ and edge set $E$, where $n=|V|$ and $m=|E|$. For a vertex $u\in V$, let $N(u)$ be its neighborhood, $N[u]=N(u)\cup\{u\}$ its closed neighborhood, and $d(u)=|N(u)|$ its degree.
@@
For a set $S\subseteq V$, its volume is $\mathrm{vol}(S)=\sum_{u\in S}d(u)$. Let $G[S]$ be the induced subgraph with vertex set $S$ and edge set $E(S)=\{\{u,v\}\in E\mid u,v\in S\}$. Let $m(S)=|E(S)|$ be the number of internal edges in $G[S]$, and for $u\in S$ define the internal degree $d_{S}(u)=|N(u)\cap S|$. The edge density of $G[S]$ is
$\eta(G[S])=\begin{cases}\dfrac{2m(S)}{|S|(|S|-1)},&|S|\geq 2,\\ 0,&\text{otherwise}.\end{cases}$
@@
Given $G=(V,E)$ and a density threshold $\gamma\in(0,1]$, a vertex set $S\subseteq V$ induces a $\gamma$-quasi-clique if the induced subgraph $G[S]$ has edge density at least $\gamma$, i.e., $\eta(G[S])\;\geq\;\gamma.$ For brevity, we also refer to $G[S]$ itself as a $\gamma$-quasi-clique.
@@
The induced subgraph $G[S^{*}]$ is a maximum $\gamma$-quasi-clique. When $\gamma=1$, MQCP reduces to the classical Maximum Clique Problem, and for any fixed $\gamma$ the problem is NP-hard and may admit multiple optimal solutions (Pattillo et al., 2013).
@@
Conductance measures how well a vertex set is separated from the rest of the graph and is widely used in local clustering (Andersen et al., 2006). For a non-empty, proper subset of vertices $S\subset V$, the cut between $S$ and its complement is defined as $\mathrm{cut}(S,V\setminus S)=|\{\{u,v\}\in E\mid u\in S,v\in V\setminus S\}|$. The conductance of $S$ is then:
@@
A smaller $\Phi(S)$ indicates better separation.
3. The EDQC Framework
This section presents EDQC for MQCP under an explicit density threshold $\gamma$. We use EDQC to denote the proposed framework. Algorithm 1 provides an instantiation of this framework (referred to as the EDQC algorithm when discussing pseudocode and complexity).
EDQC iterates over vertices as diffusion sources in non-increasing order of degree and maintains the best feasible solution found across all processed sources (lines 2–6). This ordering follows a standard efficiency heuristic (Rossi et al., 2014) that prioritizes high-degree sources, which tend to yield larger candidate regions early and thus increase the chance of finding a large feasible quasi-clique in fewer iterations. For each source vertex $v$, EDQC first computes an energy score $f:V\to\mathbb{R}_{\geq 0}$ via an adaptive diffusion procedure, providing a locality-aware ranking of vertices around $v$ (line 3). It then converts this ranking into a feasible quasi-clique by extracting a candidate set and performing density-constrained refinement to ensure the output satisfies the target threshold $\gamma$ (line 4). After computing $S$ for the current source, EDQC keeps it if it improves the best solution so far and returns the best $S^{*}$ at the end (lines 5–7).
@@ 6     $S^{*}\leftarrow S$;
3.1. Adaptive Energy Diffusion
The procedure (Algorithm 2) diffuses a unit amount of energy from a source vertex $v$ for $T$ synchronous rounds. It outputs an energy function $f:V\to\mathbb{R}_{\geq 0}$ that induces a locality-aware ranking around $v$. The ranking is used for candidate localization, while the density constraint $\gamma$ is enforced later by extraction and refinement.
@@ 17 $f\leftarrow f^{\prime}$;
Design rationale.
EDQC runs a fixed-step local diffusion for $T$ rounds and uses two simple mechanisms. A vertex is active if its current energy exceeds $\theta$, and the active set is $A=\{x\in V\mid f(x)>\theta\}$ (line 3). First, the released amount from an active vertex is adaptive: for $u\in A$, the algorithm releases $\delta(u)=\alpha(u)f(u)$, where $\alpha(u)$ increases with the total energy in the neighborhood of $u$ (line 7–8). This adaptivity encourages faster energy exchange when the diffusion remains inside a well-connected region around the source and slows down propagation when the mass reaches low-energy surroundings, which is consistent with diffusion-based local clustering using short-range propagation (Andersen et al., 2006; Kloster and Gleich, 2014; Chung, 2007). Second, in each round, EDQC updates only active vertices (line 5), which keeps the computation local and avoids spending time on vertices with negligible energy.
Compared to classical diffusion models such as PageRank (Page et al., 1999) and the heat kernel (Chung, 2007), EDQC does not aim to approximate a stationary distribution. Instead, it runs for a fixed $T$ rounds and uses the scores to induce a local ranking around the source.
Procedure.
Algorithm 2 initializes $f(v)=1$ and $f(x)=0$ for $x\neq v$ (line 1). In each round, it forms the active set $A$ (line 3) and uses a temporary copy $f^{\prime}$ to ensure synchronicity across updates (line 4). For each active vertex $u$ with $d(u)>0$, it computes the adaptive fraction $\alpha(u)$ and the released amount $\delta(u)=\alpha(u)f(u)$ (lines 5–8). It then stochastically allocates $\delta(u)$ to neighbors by sampling i.i.d. exponential variables and normalizing them into weights $\phi_{uw}$ (line 9–10), and updates $f^{\prime}$ for $u$ and its neighbors accordingly (line 11–13). Vertices with $d(u)=0$ have no neighbors to receive energy and are therefore skipped, leaving their energy unchanged. Each round ends by setting $f\leftarrow f^{\prime}$ (line 14), and after $T$ rounds the algorithm returns the final energy function $f$ (line 15).
@@
We summarize several basic properties of the diffusion procedure.
@@
Consider a fixed round and an arbitrary active vertex $u$ where $d(u)>0$. The synchronous updates associated with $u$ result in a change of $\delta(u)$ for $u$ and an addition of $\delta(u)\phi_{uw}$ for each neighbor $w\in N(u)$, satisfying $\sum_{w\in N(u)}\phi_{uw}=1$. Therefore, $\delta(u)=\sum_{w\in N(u)}\delta(u)\phi_{uw}$. Summing over all active vertices leads to the conclusion that $\sum_{x\in V}f^{\prime}(x)=\sum_{x\in V}f(x)$. Vertices that are not updated in this round (including those with $f(u)\leq\theta$ or $d(u)=0$) remain unchanged and thus do not affect the sum. ∎
The conservation of energy is a crucial property for the algorithm’s stability. It guarantees that the process only redistributes the initial energy, rather than arbitrarily creating or losing it.
@@
The adaptive nature of the diffusion is formalized by this proposition. The bounds on $\alpha(u)$ imply that each active vertex always releases at least half, but strictly less than all, of its current energy. Thus, the process continues to make progress in every round without becoming unstable. Moreover, the monotonic dependence of $\alpha(u)$ on the neighbors’ total energy means that vertices in high-energy neighborhoods release a larger fraction of energy, so diffusion proceeds more rapidly within dense regions and more slowly outside these dense regions.
@@
The expected fraction of energy that an active vertex $u$ transfers to a specific neighbor $w\in N(u)$ is uniform across all neighbors, that is, $\mathbb{E}[\phi_{uw}]=1/d(u)$.
@@
Conditioned on the current energy $f$, for an active vertex $u$ and any neighbor $w\in N(u)$, the expected update satisfies $\mathbb{E}[f^{\prime}(w)]=f(w)+\delta(u)/d(u)$. Thus, the diffusion is uniform over neighbors in expectation; any energy concentration arises from repeated exchanges along the graph structure (as in short-range diffusion for local clustering (Andersen et al., 2006; Kloster and Gleich, 2014; Chung, 2007)).
@@
EnergyDiffusion runs in $O(T(n+m))$ time and uses $O(n+m)$ space.
@@
The algorithm executes $T$ rounds. In each round, identifying the active set $A$ incurs a linear scan cost of $O(n)$. The subsequent diffusion updates iterate over active vertices, and for each $u\in A$ the work is proportional to $d(u)$. Thus the total update cost in one round is $O(\sum_{u\in A}d(u))\leq O(\sum_{u\in V}d(u))=O(m)$. Consequently, each round takes $O(n+m)$ time, yielding a total time complexity of $O(T(n+m))$. The space usage is $O(n+m)$, dominated by the graph storage and the two auxiliary energy functions $f$ and $f^{\prime}$. ∎
3.2. Quasi-Clique Extraction
The procedure (Algorithm 3) converts the energy function $f$ into a $\gamma$-quasi-clique by first ranking active vertices by energy, then selecting an initial candidate via a conductance sweep cut over energy-ranked prefixes, and finally refining the candidate to satisfy the density threshold $\gamma$. Specifically, the sweep cut selects the minimum-conductance prefix to turn the ranking into an explicit candidate region. The refinement step then enforces feasibility under $\gamma$ through pruning and greedy expansion.